Can someone please help me come up with the proper inequality equation? Thanks!
Professor Johnson gives only a midterm exam and a final exam. The semester average is computed by taking 1/3 of the midterm exam score plus 2/3 of the final exam score. The grades are determined from the semester averages by using the table below. If Stan scored only a 64 on the midterm, then for what range of scores on the final exam will Stan get a C or better? Round your answer to one decimal place.
Grading Scale
90 − 100 A
80 − 89 B
70 − 79 C
60 − 69 D

Answer by mathslover(132)

Let stan's marks in the end sem be x
therefore his average marks = 1/3 * 64 + 2/3 * x
in order for Stan to get a grade C or greater
he average marks should be greater than or equal to 70
therefore we have the inequality
1/3* 64 + 2/3 * x >= 70
Multiplying by 3 on both sides of the inequality
64 + 2x >= 210
2x >= 210-64
2x >= 146
x >= 73
therefore his range of marks for which he would get a C or better is
73 to 100